OpenStudy (anonymous):

evaluate the polynomial x = 0 3x^2-4x+1

OpenStudy (luigi0210):

at x=0?

OpenStudy (anonymous):

yess

OpenStudy (luigi0210):

Just plug in 0 for x I suppose: f(0)=3(0)^2-4(0)+1

OpenStudy (anonymous):

3x^2-4x+1=3\times0^2-4\times0+1=0+0+1=1

OpenStudy (luigi0210):

so yea, f(0)=1

OpenStudy (anonymous):

okay